OpenStudy (skullpatrol):

Why does the expression $$0^0$$ have no meaning?

OpenStudy (praxer):

https://cs.uwaterloo.ca/~alopez-o/math-faq/mathtext/node14.html This might help.........:)

OpenStudy (skullpatrol):

Thanks @praxer :)

OpenStudy (conqueror):

It's the same as 0 0's multiplied together but since it's 0 0's it has no meaning.

OpenStudy (conqueror):

I don't know if that makes any sense ^

OpenStudy (skullpatrol):

How would you explain say -2 3's then @Conqueror?

OpenStudy (conqueror):

That actually makes sense, because there is an answer to it. -2^3 = -8.

OpenStudy (anonymous):

you cant multiply anything by 0 and have it as an answer

OpenStudy (phebe):

answer choices

OpenStudy (anonymous):

if that makes sense

OpenStudy (anonymous):

that is open to debate it is not to my mind an interesting debate, but a debate nonetheless some folks say it has to be 1 most say it is undetermined

OpenStudy (anonymous):

I wonder why x^0 equals 1

OpenStudy (skullpatrol):

@satellite73 in math we don't have "debates" sir :-)

OpenStudy (anonymous):

wanna bet?

OpenStudy (skullpatrol):

Not at this level then.

OpenStudy (skullpatrol):

This is an elementary algebra question.

OpenStudy (tkhunny):

Sometimes, we can argue or not argue, but if you want something consistent with the rest of the system, you just go with it.

OpenStudy (jadeishere):

Is this a serious question because it's pretty easy to answer?

OpenStudy (cwrw238):

1 = x^ a / x^a = x^(a-a) = x^0

OpenStudy (skullpatrol):

That @cwrw238 is true for x not equal to 0.

OpenStudy (cwrw238):

yes - i was just answering BlackLabel's question

OpenStudy (anonymous):

0^0 reminds me of infinity^0

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ \bf 0^{0}=0^{1-1}=\frac{0^{\LARGE 1}}{0^{\LARGE 1}} =\LARGE\frac{0}{0} }\)

OpenStudy (skullpatrol):

but @SolomonZelman 0 has no reciprocal

OpenStudy (solomonzelman):

and ?

OpenStudy (skullpatrol):

So 0^0 has no meaning because division by 0 has no meaning.

OpenStudy (solomonzelman):

yes, that is what I am trying to say. I am not saying it is =1.

OpenStudy (cwrw238):

yes 0/0 has no meaning so 0^0 has no meaning

OpenStudy (tkhunny):

One does have to watch practical applications. There is more than one language out there that deliberately interprets 0/0 as Unity (1). It admits that it does this. It encourages the programmer to be SURE it is clear what is going on.

OpenStudy (skullpatrol):

True @tkhunny ... but given the elementary algebra definition of $$\frac{a^m}{a^n}=a^{m-n}$$ this is the best answer I have found at that level of understanding :)

OpenStudy (tkhunny):

No objection. I'm saying only that whatever argument we present, it doesn't matter a hill of beans if your programming language does something else. Good discussion.

geerky42 (geerky42):

This may help: http://youtu.be/BRRolKTlF6Q?t=6m41s

Miracrown (miracrown):

@tkhunny I agree with you absolutely. And that was coming up a bit in the discussion on it that I was reading. As a programmer myself, it is definitely something we need to be conscious of. So it sounds like your question is somewhat complex, there is a bit of discussion around the topic, with some disagreement among math teachers and mathematicians, (b/c you could show mathematically that it should be 1), unfortunately I'm not finding a great explanation of why it should be undefined, so I still can't find a convincing argument for it, actually here's the best thing I've seen . . . . so any number (non-zero) raised to the 0 power is 1, and 0 raised to any non-zero power is 0, so if we let x = 0, and follow those rules, then we get stuck with two different answers for 0^0. (the first line saying that it would be 1, the second saying it would be 0) since we can't have that conflict, many say it should be undefined. (and for what it's worth, some computer programs interpret 0^0 as 1 as well)